OpenStudy (anonymous):

Which pair of radicals is a pair of like radicals? \[9\sqrt{2} and -9\sqrt{3}\] \[7\sqrt{5} and 7\sqrt{3}\] \[5\sqrt{6} and 6\sqrt{5}\] \[3\sqrt{7} and 11\sqrt{7}\]

OpenStudy (solomonzelman):

very good with codes at the beginning, but you can use ~ for space.

OpenStudy (solomonzelman):

like radicals are the ones that can be added (saying, the once that are like terms)

OpenStudy (anonymous):

Cant they all be added?

OpenStudy (solomonzelman):

Where do you see the same radical (square root) in both terms here? In which answer choice?

OpenStudy (anonymous):

C?

OpenStudy (solomonzelman):

no, because the first one has 6 and the second one has 5, and you can't add them

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

or D?

OpenStudy (solomonzelman):

which one do you choose b or d?

OpenStudy (anonymous):

D

OpenStudy (solomonzelman):

which ones have the same square root?

OpenStudy (solomonzelman):

yes, D

OpenStudy (solomonzelman):

And in fact when you add them you get \(\large\color{black}{ \displaystyle 3\sqrt{7}+11\sqrt{7} =14\sqrt{7}}\)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

oh. Nvm

OpenStudy (anonymous):

My math was wrong